OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + y = 5 2x + y = 7 -x - 5 = 7 - 2x 5 - x = 7 - 2x 5 - x = 2x - 7

OpenStudy (anonymous):

5 - x = 7 - 2x

OpenStudy (anonymous):

x + y = 5 solve for y here y=5-x 2x + y = 7 solve for y,, y=7-2x then y=y 5-x = 7-2x